OpenStudy (anonymous):

Find a number whose product with 6 is the same as its sum with 60. (Hint: you may want to try guess and check or a table.) Be sure to check your answer for reasonableness.

OpenStudy (anonymous):

Do you understand how to turn word problems into equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

we are looking for a number in the end it'll be 6, or -3 or even 0 we just don't know yet so, until we do we'll call it x

OpenStudy (anonymous):

OK with that idea?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We know something about x we know that if we add 6 to it, we will have x + 6

OpenStudy (anonymous):

sorry add 60

OpenStudy (anonymous):

Do you understand that?

OpenStudy (anonymous):

o ok so it would be 60x

OpenStudy (anonymous):

add 60 will give x + 60 multiply by 6 will give 6x

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

really OK? The last fact we're given is that if add 60 to get x + 60 and separately, multiply by 6 to 6x those two new numbers are equal! x + 60 = 6x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you know how to work with that equation so that it will be come x = something

OpenStudy (anonymous):

no i need help with that too

OpenStudy (anonymous):

The basic rule for equations is simple: do the same thing to both sides and you still have a valid equation. x + 60 = 6x what should we do?

OpenStudy (anonymous):

subtract 6 from both sides

OpenStudy (anonymous):

x + 60 - 6 = 6x - 6 x + 54 = 6x - 6 is it getting simpler?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Nooooooo

OpenStudy (anonymous):

Try subtracting x from both sides. After all, it's just a number

OpenStudy (anonymous):

6+54=6

OpenStudy (anonymous):

Back to the original, I mean: x + 60 = 6x x + 60 - x = 6x - x 60 = 5x is that simpler?

OpenStudy (anonymous):

yea i quess

OpenStudy (anonymous):

It looks simpler to me. What we did is isolating x on one side. Next we'll multiply both sides by something. What should it be?

OpenStudy (anonymous):

60 = 5x 60 (1/5) = 5x (1/5) = x x = 12 Check your work. Is 12 + 60 = 6(12) ???

OpenStudy (anonymous):

Gotta go. HTH